OpenStudy (anonymous):

Evaluate the limit. lim x--->0 sqrt(7-x) - sqrt(7+x)/x I got 0 for the limit and my textbook says its wrong.

OpenStudy (anonymous):

multiply by the conjugate

OpenStudy (anonymous):

i did and i got - x/ 2sqrt(7)

OpenStudy (anonymous):

you forgot you had an \(x\) in the denominator

OpenStudy (anonymous):

yeah, but i canceled the -2x in the numerator by the x in the denominator...

OpenStudy (anonymous):

when you cancel and \(x\) with a\(-2x\) you are left with \(-2\) not \(-x\)

OpenStudy (anonymous):

do i multiply the numerator and denominator by sqrt(7-x) + sqrt(7+x)/sqrt(7-x) + sqrt(7+x)?

OpenStudy (anonymous):

my textbook says the limit is -1/sqrt(7). So i am not sure what i did wrong :/

OpenStudy (anonymous):

oh, nvm, i think i see my problem :)

OpenStudy (anonymous):

actually, im still confused

OpenStudy (anonymous):

should get \[\frac{-2}{\sqrt{7-x}+\sqrt{7+x}}\] then put \(x=0\)

OpenStudy (anonymous):

ah, i see now. the 2 cancels out

OpenStudy (anonymous):

you get \[\frac{-2}{\sqrt{7}+\sqrt{7}}=\frac{-1}{\sqrt7}\]

OpenStudy (anonymous):

yes, lot of cancelling here

OpenStudy (anonymous):

ok, thanks for your help :)

OpenStudy (anonymous):

yw